OBSERVATIONS AND ANALYSIS OF EXTENDED TAIL TOWARD RED IN THE DIFFUSE INTERSTEL-LAR BANDS OF HERSCHEL 36

\underline{TAKESHI OKA}, ; DANIEL E. WELTY, SEAN JOHNSON, DONALD G.YORK, LEW M. HOBBS, Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637; and JULIE DAHLSTROM, Department of Physics and Astronomy, Carthage College, Kenosha, WI 53140.

In the studies of the Diffuse Interstellar Bands (DIBs), the sightline toward Herschel 36 near the center of the HII region Messier 8 is unique. It shows spectra of CH^{+} and CH in the first excited level indicating the presence of a cloud with high radiative temperature. The heating is most likely due to far infrared emission from the adjacent intense infrared source Her 36 SE at a distance of 0.25” from Her 36.^{a}

The effect of the high radiative temperature on some DIBs is spectacular. It produces on a normally symmetric bell-shape line a very prominent Extended Tail toward Red (ETR) on prototypical DIBs λ 5780.5, λ 5797.1, and λ 6613 while other DIBs λ 5849.8, λ 6196.0, and λ 6379.3 are little affected. We interpret this as indicating that the carriers of the former 3 DIBs that are seriously affected by the radiation are polar molecules and the pronounced ETRs are the result of the decrease of rotational constant B (3 - 5 %) upon electronic excitation. High J rotational levels are pumped radiatively and with the negative (B’ - B) produces the ETR.

We have developed a model calculation of rotational distribution taking into account of both radiative and collisional processes. In view of the complexity of the problem linear molecules are considered. 7 parameters enter into the calculation but we find the fractional variation of B and the radiative temperature T_{r} are the most decisive. Although molecules with a general shape is beyond the scope of this work, we conclude that the 3 DIBs which show the pronounced ETRs are due to polar molecules and the requirement of high variation of B indicates that the molecules are not that large perhaps composed of 3-6 heavy atoms. The 3 DIBs that do not show the pronounced ETRs are likely due to non-polar molecules or large polar molecules with small fractional variation of B.^{b}

aGoto, M., Stecklum, B., Linz, H., Feldt, M., Henning, Th., Pascucci, I., and Usuda, T. 2006, ApJ, 649 299.

bOka, T., Welty, D. E., Johnson, S., York, D. G., Dahlstrom, J., and Hobbs, L. M. 2013, ApJ, submitted